HUBBLE SPACE TELESCOPE DAILY REPORT #1189 PERIOD COVERED: 0700 EST Note: Starting tomorrow the Daily Report period will cover from 0000Z (i.e., UTC) through 0000Z the following day. This change is to bring the Daily Report in line with the reporting period used on other HST reports. Daily Status Report as of 208/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1/2 5328 (HS1700+6416: A Laboratory for Astrophysical EUV Spectroscopy: Cycle 4) The High Resolution Spectrograph (Both Sides 1 and 2) was used to make ultraviolet observations of HS1700+641 as part of a study for the Laboratory for Astrophysical Research. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed FOS/RD 5532 (FOS Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to monitor the instrument's background (dark) images. The observations were completed as scheduled, and no problems were reported. 1.3 Completed WF/PC-2 5568 (WF/PC-2 Cycle 4 Calibration: DECON) The WF/PC-2 was used to make internal flat, K-spots, and dark calibration observations as part of a decontamination operation. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed FOS/RD 5642 (Observations of the Impact of the Comet Shoemaker-Levy (1993E) and Jupiter) The Faint Object Spectrograph (Red Side) made observations of the Io torus to support the Jupiter/ Shoemaker-Levy event. The observations were completed as scheduled, and no problems were reported. 1.5 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make visual flats calibration observations. The observations were executed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled :6 Successful :6 2.3 Significant HSTARs: None 2.4 Operations Notes: For the current twenty-four hour reporting period, the HST observatory has been operating in a nominal fashion. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-